a parashutist jumps from height 100 m. He wants to reach at ground with zero velocity. For this purpose he switches on a parachute propeller after falling freely for certain height. Given that after the parachute propeller is switched on total acceleration of the man varies with velocity as a=-2v, where v is the instantaneous velocity of the man . Find the time after  falling freely man should switch on parachute propeller for this purpose.

Grade:12

1 Answers

Jishnu raj paraShar
7 years ago
 
 
a=-2v
=> dv=-2dx
=>v=2x
=>4x^2 =2000-20x
=>x^2 + 5x -500=0
Solving , x=20
100-20=80
Using 2nd equation of motion,
t = 4 s